OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1 a. x = 0 b. x = 2 c. There is no x value as there is no solution. d. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

okay so basically you have to set on of the equations equal to y and then plug that value you get for y into the other equation

OpenStudy (anonymous):

So solve one of them them put the answer in the other equation?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

How would you solve the equation?

OpenStudy (lynfran):

eq.1 2x+y=1 eq.2 4x+2y=-1 from eq.1 we solve for y so y=1-2x now sub y=1-2x in eq.2 so, 4x+2(1-2x)=-1

OpenStudy (anonymous):

okay lets pic the second equation 4x + 2y = -1 subtract the 4x 2y = -1 - 4x divide by 2 y = (-1 - 4x)/2

OpenStudy (anonymous):

if you use the first equation the x's cancel out so i didn't use it

OpenStudy (anonymous):

so now plug in y = (-1 - 4x)/2 as the y for your fist equation 2x+y = 1

OpenStudy (anonymous):

so it would be 2x + (-1-4x)/2=1 ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and know just solve for x

OpenStudy (anonymous):

however im starting to think its d

OpenStudy (lynfran):

exactly the x cancels out which is true so there is an option for that...it will still cancel out ussing your way @Jdosio

OpenStudy (anonymous):

Thank you for the help it makes more sense how to solve it now.

OpenStudy (anonymous):

oh really... thats strange im getting -3/2 is it canceling out for you ?

OpenStudy (anonymous):

Yeah it is

OpenStudy (anonymous):

So wouldnt it be d if it was canceling out?

OpenStudy (anonymous):

anyway depending on how you decide to go about it you get different values for x which leads me to believe that it is d

OpenStudy (lynfran):

i would like to think its C"

OpenStudy (anonymous):

Why c?

OpenStudy (lynfran):

because the x value cancels leaving -1/2=1 ...

OpenStudy (anonymous):

But if the values get canceled out then wouldnt that mean there was just more then one solution

OpenStudy (anonymous):

it was my understanding that if you could manage to get any x value then you could cancel out c

OpenStudy (anonymous):

i place my bets on d :p

OpenStudy (anonymous):

But the equation wasnt set up like that

OpenStudy (lynfran):

but let me know if it was D...